OpenStudy (anonymous):

slope 3/4, y intercept: 9 write an equation of the line with the given slope and y-intercept

jigglypuff314 (jigglypuff314):

slope-intercept form: y = mx + b when m = slope and b = y-intercept so plug m = 3/4 and b = 9 into the equation and that would be your equation :D

OpenStudy (anonymous):

@jigglypuff314 told you what to do, If y=mx+b all you have to do is substitute the variables with the values the @jigglypuff314 has already kindly provided you with. Then put in X and work your way to Y

OpenStudy (anonymous):

so will the answer be 9=3/4=b and when write it in a equation its y=-2x=11 @Gin_Ichimaru

OpenStudy (anonymous):

Y=mx+b Y=3/4+9 I am going to use four as X Y=3/4+9 Y=3+9 Y=12

jigglypuff314 (jigglypuff314):

y = (3/4)x + (9) is your equation because m = 3/4 and b = 9 for y = mx + b

OpenStudy (anonymous):

but how would i graph that ? @jigglypuff314

jigglypuff314 (jigglypuff314):